TOIMII. (10 points) Compute V. for the sections shown. Use fc = 4,000 psi and fy= 50,000 psi. 5 in #3 stirrups 8 in. 27 in.
Answer #1

1. Given data fo= 40oopsi fyr = 50,000psi b = 18 in d = 27in Au = 3 bar= 0.22int. spacing s= 8in we know that Vn=Vctus. eunsFor I- section b = 16 in bw asin. Av = 4 = 0.goint (single legged] [Stirrups I sa bin d = 32-2.5 = 29.5in. dun d (vc+Vs) Ve=Problem 2: fyt = 60,000 pse fé = 4ooopsi Av & u urstirrups . a 0.20 inry2 = orvoine b = 150. d = 25.5infactored load Wo = 1.2x 4 = u.8Klft PL = 1.6x20= 32kleach) factored shear force at each end (V) 18] +32+32 =752k. 2 SF @ B =AAA 32k 32k WD=4.8kift mmmm B s & 6 7 75-2k. Ub.uk fuu (-) Tuink 46 -- SFD 75 . so, Vu= 75.2k. we know that dvna Vu. Vn = VuNss Ax fyd 0-40x 60.000 x 25.5 51.88x103 - 0 uox 60,000x25-5 S= 11.796 in u llin. manimum spacing = dlz (orozain - 255 (org s